# TPP-SD: Accelerating Transformer Point Process Sampling with Speculative Decoding

### v1

@@ 2.1 Temporal Point Processes
where $G(T|\mathcal{H}_{t_{N}})=\int_{0}^{T-t_{N}}g(\tau|\mathcal{H}_{t_{N}})\,\mathrm{d}\tau$. Eq. 1 is CIF-based, and Eq. 2 is CDF-based.
@@ 4 Methodology
As introduced in Sections 2.2 and 2.3, the thinning algorithm and SD share significant structural similarities despite their different application domains. Both methods adopt a two-stage “propose-verify” framework that uses a draft model to generate candidates and uses the target model for verification. Besides, the efficiency of both methods critically depends on the alignment between the proposal and target distributions.
@@
Figure 1: (a) Overall architecture of our proposed CDF-based Transformer TPP. (b) The visualization of our proposed TPP-SD sampling method as elaborated in Section 4.3.
@@ 4.2 CDF-based Transformer TPPs
Our proposed CDF-based Transformer TPP model is illustrated in Fig. 1(a). This model can be formally represented as =$\mathcal{M}=\{\mathcal{E},g(\tau|\cdot),f(k|\cdot)\}$, where $\mathcal{E}$ corresponds to the encoder and $g(\tau|\cdot),f(k|\cdot)$ correspond to the decoder. The model is trained by maximizing the log-likelihood in Eq. 2.
@@
The next timestamp is easy to sample from the log-normal mixture distribution $g_{\boldsymbol{\theta}}(\tau_{i+1}|\mathbf{h}(t_{i}))$ by (see Section A.1 for proof)
@@ 4.3 TPP-SD
Suppose we have a trained target Transformer TPP model $\mathcal{M}_{T}$ with a large number of parameters, from which we wish to sample events, and a trained draft Transformer TPP model $\mathcal{M}_{D}$ with fewer parameters, which we use to approximate $\mathcal{M}_{T}$. The procedure of TPP-SD consists of three steps, as shown in Fig. 1(b) and Algorithm 1.
@@
where $\text{norm}(\cdot)$ denotes the normalization. The correctness and necessity of sampling from the adjusted distribution have been established in the original speculative decoding work for LLMs (leviathan2023fast, ). In this work, we extend the proof to the TPP domain, as presented in Section A.2.
@@
Accept $\hat{\tau}_{i+l}$ if $\epsilon<\alpha$, where $\epsilon\sim\text{Uniform}(0,1)$. This acceptance-rejection procedure generates samples from the adjusted distribution $g^{\prime}(\tau_{i+l}|\cdot)$ defined in Eq. 3.
@@ 5.1 Evaluation of Sampling Quality and Speed
For synthetic data, where the ground truth is known, we generate samples using both AR sampling and TPP-SD, and measure the discrepancy between the ground-truth likelihood (Eq. 1) and the model likelihood of the generated samples (Eq. 2). Specifically, we compute $\Delta\mathcal{L}_{\text{ar}}^{\text{syn}}=|\mathcal{L}_{\text{gt}}-\mathcal{L}_{\text{ar}}|$ for AR sampling and $\Delta\mathcal{L}_{\text{sd}}^{\text{syn}}=|\mathcal{L}_{\text{gt}}-\mathcal{L}_{\text{sd}}|$ for TPP-SD. For real data, where the ground truth is unknown, we measure the discrepancy between the AR sampling likelihood and the TPP-SD likelihood, $\Delta\mathcal{L}^{\text{real}}=|\mathcal{L}_{\text{ar}}-\mathcal{L}_{\text{sd}}|$. Ideally, samples generated by TPP-SD and AR sampling should follow the same distribution—i.e., the ground-truth distribution. A lower likelihood discrepancy indicates higher sample quality.
@@
For real data where the ground truth is unknown, the KS statistic cannot be computed. Instead, we assess the sampling consistency between AR sampling and TPP-SD using the Wasserstein distance $D_{\text{WS}}$. Specifically, using the first $M$ events as history, we perform $N$ independent repetitions of sampling $(M+1)$-th event, yielding $\{(t_{i}^{\text{AR}},k_{i}^{\text{AR}})\}_{i=1}^{N}$ from AR sampling and $\{(t_{i}^{\text{SD}},k_{i}^{\text{SD}})\}_{i=1}^{N}$ from TPP-SD. For the temporal distribution, we compute $D_{\text{WS}}^{t}$, the 1-Wasserstein distance between the empirical distributions of $\{t_{i}^{\text{AR}}\}_{i=1}^{N}$ and $\{t_{i}^{\text{SD}}\}_{i=1}^{N}$, using ot.wasserstein_1d from the POT library (flamary2021pot, ). For the event type distribution, we compute $D_{\text{WS}}^{k}$, the earth mover’s distance between the empirical distributions of $\{k_{i}^{\text{AR}}\}_{i=1}^{N}$ and $\{k_{i}^{\text{SD}}\}_{i=1}^{N}$, using ot.emd2 from the same library. A smaller $D_{\text{WS}}$ indicates higher sample quality.
@@ 5.2 Experimental Results on Synthetic Data
We consider three synthetic datasets: inhomogeneous Poisson, univariate Hawkes, and multivariate Hawkes processes, each with 1000 sequences within the time window $[0,100]$. For each dataset, we train an 8-head, 20-layer target model and a 1-head, 1-layer draft model. We compare two sampling approaches: AR sampling using only the target model versus TPP-SD which combines both target and draft models as elaborated in Section 4.3. Details on data simulation procedures, data splitting, and experimental settings are provided in Sections B.1 and C.3.1.
@@ Results.
As shown in Table 1, across all three synthetic datasets and three encoder architectures, TPP-SD consistently demonstrates high-fidelity sampling that closely matches the performance of AR sampling. Specifically, both TPP-SD and AR sampling exhibit low likelihood discrepancies and near-zero KS statistics relative to the ground truth, indicating excellent alignment with the target distribution. Furthermore, as shown in Fig. 2, the KS plots show that the samples generated by both methods consistently fall within the 95% confidence bands.
@@
Wall-time $T$ ($\downarrow$) AR Sampling 3.477 2.680 12.103 5.147 2.747 20.503 4.007 2.490 12.403
@@
In addition, architectural variations are evident: AttNHP achieves the highest acceleration ratios despite having the slowest wall times; SAHP shows the lowest speedup but maintains the fastest wall times; and THP strikes a balance between acceleration ratio and runtime. These results suggest that the underlying architecture plays a significant role in determining the effectiveness of acceleration. A detailed discussion of these differences is provided in Section D.2.
@@
Wall-time $T$ ($\downarrow$) AR Sampling 5.890 2.460 16.256 1.023 0.900 7.657 1.157 1.183 2.573 1.353 1.423 3.217
@@ 5.3 Experimental Results on Real Data
For real data, we consider four commonly used datasets: Taobao (tianchi-taobao-2018, ), Amazon (ni2019justifying, ), Taxi (whong2014foiking, ), and StackOverflow (nan2016recurrent, ). We similarly train an 8-head, 20-layer target model and a 1-head, 1-layer draft model as in Section 5.2, and compare the sampling performance between AR sampling and TPP-SD. Details on data statistics, data splitting, and experimental settings are provided in Section C.3.2. It is worth noting that for real data, the ground-truth distribution is unknown. Therefore, we use AR sampling as a reference and quantify how closely TPP-SD approximates it. To establish a meaningful baseline that accounts for the inherent stochasticity of TPP sampling, we compare two independent runs of AR sampling. In theory, the likelihood discrepancy and Wasserstein distances between two such runs should be zero. However, due to stochastic variation, even independent autoregressive runs may exhibit small differences. This self-comparison serves as a baseline for evaluating how well TPP-SD aligns with AR sampling.
@@ 5.4 Ablation Studies
Dataset Encoder Type Draft Model $\Delta\mathcal{L}$ Distance $\alpha$ ($\uparrow$) Wall-time Speedup Ratio
head layer $D_{\text{KS}}$ ($\downarrow$) $D_{\text{WS}}^{t}$ ($\downarrow$) $D_{\text{WS}}^{k}$ ($\downarrow$) $T_{\text{AR}}$ ($\downarrow$) $T_{\text{SD}}$ ($\downarrow$) $S_{\text{AR/SD}}$ ($\uparrow$)
@@ A.2 Correctness of Sampling from the Adjusted Distribution
Denote $P(\text{accept},\tau^{\prime})$ as the probability that the sample $\tau^{\prime}$ drawn by the draft model is accepted by the target model, and $P(\text{reject},\tau^{\prime})$ as the probability that $\tau^{\prime}$ is rejected and subsequently resampled from the adjusted distribution defined in Eq. 3. By the law of total probability, the distribution of a sample $\tau^{\prime}$ drawn by SD can be calculated as
@@ A.3 Proof of Theorem 1
which justifies the correctness of acceptance-rejection sampling. Now we apply this acceptance-rejection sampling framework to the specific scenario described in the theorem. The adjusted distribution $g^{\prime}(\tau_{i+l}|\cdot)$ defined in Eq. 3 corresponds to the target distribution $p(\tau)$, and $g_{T}(\tau_{i+l}|\cdot)$ corresponds to the proposal distribution $q(\tau)$. We need to find the constant $M$ such that for any sampled $\hat{\tau}_{i+L}$, $g^{\prime}(\hat{\tau}_{i+l}|\cdot)\leq Mg_{T}(\hat{\tau}_{i+l}|\cdot)$. Denote $Z=\int\max\left(0,g_{T}(\hat{\tau}_{i+l}|\cdot)-g_{D}(\hat{\tau}_{i+l}|\cdot)\right)\mathrm{d}\hat{\tau}_{i+l}$, note that
@@
which is precisely the acceptance threshold $\alpha$ defined in Theorem 1.
@@ A.4 Computation of the KS Statistic
We employ the Kolmogorov-Smirnov (KS) statistic to assess whether sampled realizations conform to ground truth distributions in synthetic point processes. For $n$ i.i.d. ordered observations $\{X_{i}\}_{i=1}^{n}$ from a distribution $F(x)$, the KS statistic is defined as $D_{\text{KS}}=\sup_{x}|F_{n}(x)-F(x)|$, where $F_{n}(x)$ is the empirical cumulative distribution of $\{X_{i}\}_{i=1}^{n}$. The leverage of KS statistic is guaranteed by Theorem 2.
@@ C.1 Algorithm for TPP-SD
The pseudo code for the algorithm of TPP-SD is shown in Algorithm 1.
@@ C.3.1 Sampling Synthetic TPPs
For the experiments in Section 5.2, we evaluate sampling performance within the time window $[0,100]$. We report the discrepancy metrics $\Delta\mathcal{L}_{\text{ar}}^{\text{syn}}$ and $\Delta\mathcal{L}_{\text{sd}}^{\text{syn}}$, the KS statistics $D_{\text{KS}}$, the execution time $T_{\text{AR}}$ and $T_{\text{SD}}$, and the resulting speedup ratio $S_{\text{AR/SD}}$ as defined in Section 5.1. We conducted experiments across three random seeds and report the mean for each metric to ensure result robustness.
@@
We construct direct comparison visualizations between $F_{n}(x)$ and $F(x)=1-e^{-x}$ alongside KS plots across Transformer encoders (THP, SAHP, AttNHP) with draft length $\gamma=10$ on three synthetic datasets. As evidenced in Fig. 4, $F_{n}(x)$ demonstrates strong concordance with $F(x)=1-e^{-x}$, and all points in the KS plots remain within the 95% confidence bands, providing robust statistical evidence for the distributional fidelity of our TPP-SD method.
@@ C.3.2 Sampling Real-world TPPs
For the experiments in Section 5.3, we conduct evaluations within the same time window of $[0,100]$. In accordance with Section C.3.1, we report the discrepancy metric $\Delta\mathcal{L}^{\text{real}}$, the Wasserstein distance $D_{\text{WS}}^{t}$ and $D_{\text{WS}}^{k}$, the execution time $T_{\text{AR}}$ and $T_{\text{SD}}$ for AR sampling and TPP-SD, and the resulting speedup ratio $S_{\text{AR/SD}}$ defined in Section 5.1. For computation of $D_{WS}^{t}$ and $D_{WS}^{k}$, we set $M=100,N=100$, which means that we fix the first $M=100$ events as history and perform $N=100$ independent repetitions of sampling the $101$-th event, obtaining $\{(t_{i}^{\text{AR}},k_{i}^{\text{AR}}\}_{i=1}^{100}$ from AR sampling and $\{(t_{i}^{\text{SD}},k_{i}^{\text{SD}}\}_{i=1}^{100}$ from TPP-SD, and subsequently computing $D_{\text{WS}}^{t}$ and $D_{\text{WS}}^{k}$ as elaborated in Section 5.1.
@@ C.4 Ablation Studies
We have investigated the impact of draft length $\gamma$ and draft model size on the sampling quality and sampling speed of TPP-SD with AttNHP encoder. Using the same sampling configuration and evaluation metrics as in Section 5.4, we also conducted experiments using THP and SAHP encoders.
@@
As evidenced by Table 4, increasing the draft model size with either THP or SAHP encoder preserves sampling quality (measured by $\Delta\mathcal{L}$ and $D$) and improves the acceptance rate $\alpha$, but reduces the speedup ratio $S_{\text{AR/SD}}$. The 1-head, 1-layer configuration remains optimal for the draft model across these encoder architectures.
@@
Dataset Encoder Type Draft Model $\Delta\mathcal{L}$ Distance $\alpha$ ($\uparrow$) Wall-time Speedup Ratio
head layer $D_{\text{KS}}$ ($\downarrow$) $D_{\text{WS}}^{t}$ ($\downarrow$) $D_{\text{WS}}^{k}$ ($\downarrow$) $T_{\text{AR}}$ ($\downarrow$) $T_{\text{SD}}$ ($\downarrow$) $S_{\text{AR/SD}}$ ($\uparrow$)

### v2

@@ 2.1 Temporal Point Processes
where $G(T|\mathcal{H}_{t_{N}})=\int_{0}^{T-t_{N}}g(\tau|\mathcal{H}_{t_{N}})\,\mathrm{d}\tau$. Eq.˜1 is CIF-based, and Eq.˜2 is CDF-based.
@@ 4 Methodology
As introduced in Sections˜2.2 and 2.3, the thinning algorithm and SD share significant structural similarities despite their different application domains. Both methods adopt a two-stage “propose-verify” framework that uses a draft model to generate candidates and uses the target model for verification. Besides, the efficiency of both methods critically depends on the alignment between the proposal and target distributions.
@@
Figure 1: (a) Overall architecture of our proposed CDF-based Transformer TPP. (b) The visualization of our proposed TPP-SD sampling method as elaborated in Section˜4.3.
@@ 4.2 CDF-based Transformer TPPs
Our proposed CDF-based Transformer TPP model is illustrated in Fig.˜1(a). This model can be formally represented as =$\mathcal{M}=\{\mathcal{E},g(\tau|\cdot),f(k|\cdot)\}$, where $\mathcal{E}$ corresponds to the encoder and $g(\tau|\cdot),f(k|\cdot)$ correspond to the decoder. The model is trained by maximizing the log-likelihood in Eq.˜2.
@@
The next timestamp is easy to sample from the log-normal mixture distribution $g_{\boldsymbol{\theta}}(\tau_{i+1}|\mathbf{h}(t_{i}))$ by (see Section˜A.1 for proof)
@@ 4.3 TPP-SD
Suppose we have a trained target Transformer TPP model $\mathcal{M}_{T}$ with a large number of parameters, from which we wish to sample events, and a trained draft Transformer TPP model $\mathcal{M}_{D}$ with fewer parameters, which we use to approximate $\mathcal{M}_{T}$. The procedure of TPP-SD consists of three steps, as shown in Fig.˜1(b) and Algorithm˜1.
@@
[…]
@@
Accept $\hat{\tau}_{i+l}$ if $\epsilon<\alpha$, where $\epsilon\sim\text{Uniform}(0,1)$. This acceptance-rejection procedure generates samples from the adjusted distribution $g^{\prime}(\tau_{i+l}|\cdot)$ defined in Eq.˜3.
@@ 5.1 Evaluation of Sampling Quality and Speed
For synthetic data, where the ground truth is known, we generate samples using both AR sampling and TPP-SD, and measure the discrepancy between the ground-truth likelihood (Eq.˜1) and the model likelihood of the generated samples (Eq.˜2). Specifically, we compute $\Delta\mathcal{L}_{\text{ar}}^{\text{syn}}=|\mathcal{L}_{\text{gt}}-\mathcal{L}_{\text{ar}}|$ for AR sampling and $\Delta\mathcal{L}_{\text{sd}}^{\text{syn}}=|\mathcal{L}_{\text{gt}}-\mathcal{L}_{\text{sd}}|$ for TPP-SD. For real data, where the ground truth is unknown, we measure the discrepancy between the AR sampling likelihood and the TPP-SD likelihood, $\Delta\mathcal{L}^{\text{real}}=|\mathcal{L}_{\text{ar}}-\mathcal{L}_{\text{sd}}|$. Ideally, samples generated by TPP-SD and AR sampling should follow the same distribution—i.e., the ground-truth distribution. A lower likelihood discrepancy indicates higher sample quality.
@@
[…]
@@ 5.2 Experimental Results on Synthetic Data
We consider three synthetic datasets: inhomogeneous Poisson, univariate Hawkes, and multivariate Hawkes processes, each with 1000 sequences within the time window [0,100][0,100][ 0 , 100 ]. For each dataset, we train an 8-head, 20-layer target model and a 1-head, 1-layer draft model. We compare two sampling approaches: AR sampling using only the target model versus TPP-SD which combines both target and draft models as elaborated in Section˜4.3. Details on data simulation procedures, data splitting, and experimental settings are provided in Sections˜B.1 and C.3.1.
@@ Results.
As shown in Table˜1, across all three synthetic datasets and three encoder architectures, TPP-SD consistently demonstrates high-fidelity sampling that closely matches the performance of AR sampling. Specifically, both TPP-SD and AR sampling exhibit low likelihood discrepancies and near-zero KS statistics relative to the ground truth, indicating excellent alignment with the target distribution. Furthermore, as shown in Fig.˜2, the KS plots show that the samples generated by both methods consistently fall within the 95% confidence bands.
@@
Wall-time
$T$ ($\downarrow$)
AR Sampling 3.477 2.680 12.103 5.147 2.747 20.503 4.007 2.490 12.403
@@
In addition, architectural variations are evident: AttNHP achieves the highest acceleration ratios despite having the slowest wall times; SAHP shows the lowest speedup but maintains the fastest wall times; and THP strikes a balance between acceleration ratio and runtime. These results suggest that the underlying architecture plays a significant role in determining the effectiveness of acceleration. A detailed discussion of these differences is provided in Section˜D.2.
@@
Wall-time
$T$ ($\downarrow$)
AR Sampling 5.890 2.460 16.256 1.023 0.900 7.657 1.157 1.183 2.573 1.353 1.423 3.217
@@ 5.3 Experimental Results on Real Data
For real data, we consider four commonly used datasets: Taobao (tianchi-taobao-2018, ), Amazon (ni2019justifying, ), Taxi (whong2014foiking, ), and StackOverflow (nan2016recurrent, ). We similarly train an 8-head, 20-layer target model and a 1-head, 1-layer draft model as in Section˜5.2, and compare the sampling performance between AR sampling and TPP-SD. Details on data statistics, data splitting, and experimental settings are provided in Section˜C.3.2. It is worth noting that for real data, the ground-truth distribution is unknown. Therefore, we use AR sampling as a reference and quantify how closely TPP-SD approximates it. To establish a meaningful baseline that accounts for the inherent stochasticity of TPP sampling, we compare two independent runs of AR sampling. In theory, the likelihood discrepancy and Wasserstein distances between two such runs should be zero. However, due to stochastic variation, even independent autoregressive runs may exhibit small differences. This self-comparison serves as a baseline for evaluating how well TPP-SD aligns with AR sampling.
@@ 5.4 Ablation Studies
Dataset
Encoder
Type
Draft Model $\Delta\mathcal{L}$ Distance $\alpha$ ($\uparrow$) Wall-time Speedup Ratio
head layer
$D_{\text{KS}}$ ($\downarrow$)
$D_{\text{WS}}^{t}$ ($\downarrow$)
$D_{\text{WS}}^{k}$ ($\downarrow$)
$T_{\text{AR}}$ ($\downarrow$)
$T_{\text{SD}}$ ($\downarrow$)
$S_{\text{AR/SD}}$ ($\uparrow$)
@@ A.2 Correctness of Sampling from the Adjusted Distribution
Denote $P(\text{accept},\tau^{\prime})$ as the probability that the sample $\tau^{\prime}$ drawn by the draft model is accepted by the target model, and $P(\text{reject},\tau^{\prime})$ as the probability that $\tau^{\prime}$ is rejected and subsequently resampled from the adjusted distribution defined in Eq.˜3. By the law of total probability, the distribution of a sample $\tau^{\prime}$ drawn by SD can be calculated as
@@ A.3 Proof of Theorem 1
which justifies the correctness of acceptance-rejection sampling. Now we apply this acceptance-rejection sampling framework to the specific scenario described in the theorem. The adjusted distribution $g^{\prime}(\tau_{i+l}|\cdot)$ defined in Eq.˜3 corresponds to the target distribution $p(\tau)$, and $g_{T}(\tau_{i+l}|\cdot)$ corresponds to the proposal distribution $q(\tau)$. We need to find the constant $M$ such that for any sampled $\hat{\tau}_{i+L}$, $g^{\prime}(\hat{\tau}_{i+l}|\cdot)\leq Mg_{T}(\hat{\tau}_{i+l}|\cdot)$. Denote $Z=\int\max\left(0,g_{T}(\hat{\tau}_{i+l}|\cdot)-g_{D}(\hat{\tau}_{i+l}|\cdot)\right)\mathrm{d}\hat{\tau}_{i+l}$, note that
@@
which is precisely the acceptance threshold $\alpha$ defined in Theorem˜1.
@@ A.4 Computation of the KS Statistic
We employ the Kolmogorov-Smirnov (KS) statistic to assess whether sampled realizations conform to ground truth distributions in synthetic point processes. For $n$ i.i.d. ordered observations $\{X_{i}\}_{i=1}^{n}$ from a distribution $F(x)$, the KS statistic is defined as $D_{\text{KS}}=\sup_{x}|F_{n}(x)-F(x)|$, where $F_{n}(x)$ is the empirical cumulative distribution of $\{X_{i}\}_{i=1}^{n}$. The leverage of KS statistic is guaranteed by Theorem˜2.
@@ C.1 Algorithm for TPP-SD
The pseudo code for the algorithm of TPP-SD is shown in Algorithm˜1.
@@ C.3.1 Sampling Synthetic TPPs
For the experiments in Section˜5.2, we evaluate sampling performance within the time window [0,100][0,100][ 0 , 100 ]. We report the discrepancy metrics $\Delta\mathcal{L}_{\text{ar}}^{\text{syn}}$ and $\Delta\mathcal{L}_{\text{sd}}^{\text{syn}}$, the KS statistics $D_{\text{KS}}$, the execution time $T_{\text{AR}}$ and $T_{\text{SD}}$, and the resulting speedup ratio $S_{\text{AR/SD}}$ as defined in Section˜5.1. We conducted experiments across three random seeds and report the mean for each metric to ensure result robustness.
@@
We construct direct comparison visualizations between $F_{n}(x)$ and $F(x)=1-e^{-x}$ alongside KS plots across Transformer encoders (THP, SAHP, AttNHP) with draft length $\gamma=10$ on three synthetic datasets. As evidenced in Fig.˜4, $F_{n}(x)$ demonstrates strong concordance with $F(x)=1-e^{-x}$, and all points in the KS plots remain within the 95% confidence bands, providing robust statistical evidence for the distributional fidelity of our TPP-SD method.
@@ C.3.2 Sampling Real-world TPPs
For the experiments in Section˜5.3, we conduct evaluations within the same time window of [0,100][0,100][ 0 , 100 ]. In accordance with Section˜C.3.1, we report the discrepancy metric $\Delta\mathcal{L}^{\text{real}}$, the Wasserstein distance $D_{\text{WS}}^{t}$ and $D_{\text{WS}}^{k}$, the execution time $T_{\text{AR}}$ and $T_{\text{SD}}$ for AR sampling and TPP-SD, and the resulting speedup ratio $S_{\text{AR/SD}}$ defined in Section˜5.1. For computation of $D_{WS}^{t}$ and $D_{WS}^{k}$, we set $M=100,N=100$, which means that we fix the first $M=100$ events as history and perform $N=100$ independent repetitions of sampling the 101101101-th event, obtaining $\{(t_{i}^{\text{AR}},k_{i}^{\text{AR}}\}_{i=1}^{100}$ from AR sampling and $\{(t_{i}^{\text{SD}},k_{i}^{\text{SD}}\}_{i=1}^{100}$ from TPP-SD, and subsequently computing $D_{\text{WS}}^{t}$ and $D_{\text{WS}}^{k}$ as elaborated in Section˜5.1.
@@ C.4 Ablation Studies
We have investigated the impact of draft length $\gamma$ and draft model size on the sampling quality and sampling speed of TPP-SD with AttNHP encoder. Using the same sampling configuration and evaluation metrics as in Section˜5.4, we also conducted experiments using THP and SAHP encoders.
@@
As evidenced by Table˜4, increasing the draft model size with either THP or SAHP encoder preserves sampling quality (measured by $\Delta\mathcal{L}$ and $D$) and improves the acceptance rate $\alpha$, but reduces the speedup ratio $S_{\text{AR/SD}}$. The 1-head, 1-layer configuration remains optimal for the draft model across these encoder architectures.
@@
Dataset
Encoder
Type
Draft Model $\Delta\mathcal{L}$ Distance $\alpha$ ($\uparrow$) Wall-time Speedup Ratio
head layer
$D_{\text{KS}}$ ($\downarrow$)
$D_{\text{WS}}^{t}$ ($\downarrow$)
$D_{\text{WS}}^{k}$ ($\downarrow$)
$T_{\text{AR}}$ ($\downarrow$)
$T_{\text{SD}}$ ($\downarrow$)
$S_{\text{AR/SD}}$ ($\uparrow$)